# TEXTOC: Text-driven Object-Centric Style Transfer

### v2

@@ Appendix C Limitations of TEXTOC
A noted limitation of TEXTOC is its reliance on the CLIP model’s feature space and classification capabilities. Consequently, its performance may diminish for styles or objects less represented or absent in the CLIP training dataset, such as recent art or gadgets, as illustrated in Fig. 10. This dependency highlights a potential area for future development, aiming to enhance TEXTOC’s adaptability to emerging styles and objects.
@@
Figure 10: Failure cases of TEXTOC. The numbers following the style text denote the year of the art work was created.
Appendix D Comparison to mask-guided generative models
In the realm of generative models, significant strides have been made in image generation and editing. Yet, a prevalent challenge among these models, including those guided by masks, is their propensity to inadvertently alter the structural integrity of targeted objects within an image. This often leads to a compromise in the overall image quality and authenticity. To illustrate this limitation and to position our TEXTOC model within the context of current advancements, we conduct a comparative analysis with existing mask-guided generative models [2, 38]. These models typically rely on masks to designate specific areas for image editing. Our comparison, as depicted in Fig. 11, highlights a key distinction. Unlike mask-based models, which may inadvertently distort vital details of the object, TEXTOC consistently maintains the structural integrity of the object, focusing solely on altering its style. This distinction is crucial in preserving the overall coherence and authenticity of the image. Moreover, an important advantage of TEXTOC is its autonomous capability to recognize and apply style transfer to objects relevant to the given text. This feature eliminates the cumbersome and often imprecise process of manual mask creation, streamlining the style transfer process. TEXTOC, thus, not only preserves the integrity of the objects but also offers a more efficient and user-friendly approach to style transfer compared to traditional mask-guided generative models.
Figure 11: Comparative analysis of TEXTOC with other generative models using mask input. This figure illustrates the effectiveness of our model in handling style transfer challenges where specific object masks are used to guide the process.
Appendix E User study
@@
We conducted the user study to enhance our evaluation process, the results of which are presented in Tab. 4. This study involved 50 participants, whose ages ranged from their 20s to 50s. We designed the study to assess three key aspects: stylization quality, background preservation, and content preservation, providing nine examples within each category for evaluation. We compared the stylized images generated by our model against those produced by Text2live [3], Glide [38], and Blended Diffusion [2], all of which also aim for object-centric style transfer. Our analysis indicates that our model excels in achieving a balanced object-centric style transfer, effectively maintaining the semantic integrity of the object and the background without compromise. The user study question examples illustrated in Fig. 12.
@@
Blended Diffusion 4.4 % 7.3 % 2.4 %
Glide 1.1 % 12.0 % 3.6 %
Text2live 29.3% 14.9 % 11.8 %
Ours 65.1 % 65.7 % 82.2 %
@@
Appendix F Additional object-centric style transfer results using our TEXTOC
We show additional results of our TEXTOC method, as seen in Fig. 13, Fig. 14, Fig. 15, and Fig. 16. These figures illustrate the effectiveness of our approach in various style transfer scenarios, emphasizing the versatility of TEXTOC. TMPS plays a crucial role in initiating the style transfer process. It specifically targets image areas that correspond with the input text, ensuring that the chosen style is applied seamlessly and appropriately to the relevant objects. This targeted approach results in a harmonious blend of the new style with the original image, particularly in areas corresponding to the source text. Furthermore, our method incorporates an innovative ABP loss. This component of TEXTOC is vital in maintaining the integrity of the background areas during the style transfer process. It ensures that these areas remain unaffected by the changes applied to the object of interest. This loss function is key to achieving a balanced and natural-looking result where the style changes are confined to the targeted object, while the rest of the image retains its original appearance. The results in Fig. 13, Fig. 14, Fig. 15, and Fig. 16 collectively showcase the robust customizability and adaptability of the TEXTOC method. They demonstrate its capability to handle a diverse range of styles and scenarios, effectively adapting the chosen style to the specific objects in the image as dictated by the source text, all while preserving the overall aesthetic and integrity of the background.
Figure 13: Stylization results demonstrating various “artistic” styles guided by text using our TEXTOC model.
Figure 14: Stylization results demonstrating various “color” styles guided by text using our TEXTOC model.
Figure 15: Stylization results demonstrating various “texture” styles guided by text using our TEXTOC model.

### v3

@@ Appendix E Limitations of Style-Editor
A noted limitation of Style-Editor is its reliance on the CLIP model’s feature space and classification capabilities. Consequently, its performance may diminish for styles or objects less represented or absent in the CLIP training dataset, such as recent art or gadgets, as illustrated in Fig. 14. This dependency highlights a potential area for future development, aiming to enhance Style-Editor’s adaptability to emerging styles and objects.
@@
Figure 14: Failure cases of Style-Editor. The numbers following the style text denote the year of the art work was created.
Appendix F User study
@@
We conducted two rounds of user studies to enhance our evaluation process, the results of which are presented in Tab. 6. This study involved 50 participants, whose ages ranged from their 20s to 50s. We designed the study to assess three key aspects: stylization quality, background preservation, and content preservation, providing nine examples within each category for evaluation. The evaluation included the mask-based editing models Glide [41], and Blended Diffusion [1], along with the top-4 models from Tab. 1 — FlexIT [7], LEDITS ++ [3], Null text inversion [40], and Text2LIVE [2]. In each round, we compared the stylized images produced by three competing models with those generated by our model. Our analysis indicates that our model excels in achieving a balanced object-centric style editing, effectively maintaining the semantic integrity of the object and the background without compromise. In particular, compared with mask-based methods, our model delivers more natural and stable performance. The user study question examples are illustrated in Fig. 15.
@@
Blended Diffusion [1] 4.4 % 7.3 % 2.4 %
Glide [41] 1.1 % 12.0 % 3.6 %
Text2LIVE [2] 29.3 % 14.9 % 11.8 %
FlexIT [7] 14 % 5.1 % 2.9 %
LEDITS++ [3] 13.3 % 6 % 6.2 %
Null-text inversion [40] 5.8 % 25.8 % 16 %
Ours (Avg.) 66 % 64.4 % 78.6 %
@@
Appendix G Additional quantitative results
To assess the robustness of our model, we conducted additional quantitative experiments on the top-5 models from Tab. 1, with training time evaluated following the settings in Sec. B.3. The test set was expanded by randomly selecting 50 images from the MSCOCO 2017 dataset and pairing each with 10 different style text descriptions, yielding a total of 500 stylized images for evaluation. As presented in Tab. 7, our model consistently outperforms others on the extended test set while ranking second in speed. Notably, even compared to LEDITS++ [3], the fastest model, our approach demonstrates superior performance in both foreground and background metrics. These results validate its strong generalization ability and highlight its well-balanced stylized output at a relatively fast pace.
Foreground metrics Background metrics
[…]
FlexIT [7] 0.24 7.03 0.20 4.09 0.36 0.66 0.14 21.17 59.8
LEDITS++ [3] 0.21 5.98 0.19 2.70 0.42 0.75 0.13 21.43 10.4
NTI [40] 0.20 4.64 0.16 3.02 0.32 0.74 0.12 23.46 102.3
LPM [46] 0.20 8.62 0.25 4.81 0.73 0.67 0.19 19.53 119.9
Text2LIVE [2] 0.30 3.65 0.13 1.25 0.17 0.87 0.08 25.47 412.6
Ours 0.31 3.09 0.10 0.95 0.08 0.89 0.08 26.71 44.3
Table 7: Additional quantitative comparison with 500 stylized images. $\uparrow$ indicates that higher values are better, while $\downarrow$ indicates that lower values are better.
Appendix H Additional object-centric style editing results using our Style-Editor
We show additional results of our Style-Editor method across diverse scenes, as seen in Fig. 16, Fig. 17, Fig. 18, and Fig. 19. These figures illustrate the effectiveness of our approach in various style editing scenarios, emphasizing the versatility of Style-Editor. TMPS plays a crucial role in initiating the style editing process. It specifically targets image areas that correspond with the input text, ensuring that the chosen style is applied seamlessly and appropriately to the relevant objects. This targeted approach results in a harmonious blend of the new style with the original image, particularly in areas corresponding to the source text. Furthermore, our method incorporates an innovative ABP loss. This component of Style-Editor is vital in maintaining the integrity of the background areas during the style editing process. It ensures that these areas remain unaffected by the changes applied to the object of interest. This loss function is key to achieving a balanced and natural-looking result where the style changes are confined to the targeted object, while the rest of the image retains its original appearance. The results in Fig. 16, Fig. 17, Fig. 18, and Fig. 19 collectively showcase the robust customizability and adaptability of the Style-Editor method. They demonstrate its capability to handle a diverse range of styles and scenarios, effectively adapting the chosen style to the specific objects in the image as dictated by the source text, all while preserving the overall aesthetic and integrity of the background.
Figure 16: Stylization results demonstrating various “artistic" styles guided by text using our Style-Editor model.
Figure 17: Stylization results demonstrating various “color" styles guided by text using our Style-Editor model.
Figure 18: Stylization results demonstrating various “texture" styles guided by text using our Style-Editor model.